OpenStudy (anonymous):

Simplify (-k + 4) + (-2 + 3k).

OpenStudy (anonymous):

(-k + 4) + (-2 + 3k) -k +4 -2 +3k 2 + 2k

OpenStudy (anonymous):

Just solve what's inside the parenthesis and then combine like terms.

OpenStudy (cj49):

k=-1

OpenStudy (anonymous):

oh ok thanks

OpenStudy (mokeira):

\[(-k+4)+(-2+3k)\] open up the brackets \[-k+4-2+3k\] \[3k-k+4-2\] \[2k+2\]

OpenStudy (mokeira):

@cj49 that is not true

OpenStudy (idealist10):

(-k+4)+(-2+3k) -k+3k+4+(-2) 2k+4-2 2k+2 Answer: 2k+2

OpenStudy (cj49):

2k+2=0 k=-2/2 k= -1

OpenStudy (anonymous):

what about this question guys I want to know how you did it Simplify the following expression: 5x + 6 - x + 12

OpenStudy (anonymous):

2 + 2k = 0 Subtract 2 from both sides. 2k = -2 Divide both sides by 2. k = -1 So that is true.

OpenStudy (anonymous):

Wow I type way too slow.

OpenStudy (anonymous):

it looks like the answer was 2k+2

OpenStudy (anonymous):

Well it is 2k+2 because the problem told you to simplify.

OpenStudy (idealist10):

5x-x+6+12 =4x+18 Answer: 4x+18

OpenStudy (anonymous):

It would be k=-1 if it asked you to solve for k, but it did not.

OpenStudy (anonymous):

what about this question guys I want to know how you did it Simplify the following expression: 5x + 6 - x + 12

OpenStudy (mokeira):

it is an expression not equation...there is no equal sign

OpenStudy (mokeira):

@cj49

OpenStudy (cj49):

ohk.. i thought we have to find k

OpenStudy (mokeira):

i also equated it but when i looked again i realized it is an expression